Since there is a direct relationship between absorbance and concentration of Beer's law, the units for the slope (product of molar absorptivity and path length) are:

Question 4 options: (I chose number 1 because the molar absorvity of a solute is in units of M^-1cm^-1. Is this correct?)

1) M-1cm-1
2) M-1cm
3) M-1
4) cm-1
5)

unitless
Answer #1

A = e*l*C

we need to find the unit of e*l

e*l = A/C

A is absorbance which is unit less

unit of C is M

So, unit of e*l = 1/M = M-1

Answer: option 3